RECENT DEVELOPMENTS IN THE STEREOLOGICAL ANALYSIS OF PARTICLES

EVA B. VEDEL JENSEN

Department of Theoretical Statistics, Institute of Mathematics, University of Aarhus, Ny Munkegade, DK-8000 Aarhus C, Denmark

(Received November 14, 1989; revised March 26, 1991)

Abstract. Recent developments in the stereological analysis of particles are reviewed. The trend has been towards methods which are applicable without specific assumptions about particle shape. Geometric samples of a local 3-d character are used. Stereological estimators of particle intensity, particle size distribution and particle interaction are presented and discussed.

Key words and phrases: Intensity, K-function, mark distribution, marked point processes, measure decomposition, nucleator, second-order properties, stereology, stochastic geometry.